In a common test for cardiac function (the “stress test”), the patient walks on an

inclined treadmill. Estimate the power required from a 75-kg patient when the

treadmill is sloping at an angle of 12° and the velocity is 3.1 km/h.


1
Expert's answer
2022-05-30T15:17:25-0400
"P=Fv=mgv\sin\theta \\
=75*9.8*(3.1\div3.6)*\sin12^\circ=132\:\rm W"